OpenStudy (samara8954):

Solve for n. n = 270 n = 10 n = 50 n = 5

OpenStudy (samara8954):

@dtuniyants

OpenStudy (anonymous):

reduce 5/30=5(5*6)=1/6 n/30=1/6 n*6=30----->n=5

OpenStudy (samara8954):

Solve for n. n = 105 n = 3 n = 8 n = 24